I cannot figure out how to do this word problem:
Write an equation and solve this problem. Find the sum of three consecutive odd integers if the sum of the first two integers is equal to twenty-four less than four times the third integer.

Answer by lwsshak3(11628)

Write an equation and solve this problem. Find the sum of three consecutive odd integers if the sum of the first two integers is equal to twenty-four less than four times the third integer.
**
let x=first consecutive odd integer
x+2=second consecutive odd integer
x+4=third consecutive odd integer
..
x+(x+2)=4(x+4)-24
x+x+2=4x+16-24
2x-4x=16-24-2
-2x=-10
x=-10/-2=5
ans:
first consecutive odd integer=5
second consecutive odd integer=x+2=7
third consecutive odd integer=+4=9